# Dynamic Parameter Memory: Temporary LoRA-Enhanced LLM for Long-Sequence Emotion Recognition in Conversation

### v1

@@ Abstract
Recent research has focused on applying speech large language model (SLLM) to improve speech emotion recognition (SER). However, the inherently high frame rate in speech modality severely limits the signal processing and understanding capabilities of SLLM. For example, a SLLM with a 4K context window can only process 80 seconds of audio at 50Hz feature sampling rate before reaching its capacity limit. Input token compression methods used in SLLM overlook the continuity and inertia of emotions across multiple conversation turns. This paper proposes a Dynamic Parameter Memory (DPM) mechanism with contextual semantics and sentence-level emotion encoding, enabling processing of unlimited-length audio with limited context windows in SLLM. Specifically, DPM progressively encodes sentence-level information and emotions into a temporary LoRA module during inference to effectively "memorize" the contextual information. We trained an emotion SLLM as a backbone and incorporated our DPM into inference for emotion recognition in conversation (ERC). Experimental results on the IEMOCAP dataset show that DPM significantly improves the emotion recognition capabilities of SLLM when processing long audio sequences, achieving state-of-the-art performance.
Dynamic Parameter Memory: Temporary LoRA-Enhanced LLM for Long-Sequence Emotion Recognition in Conversation
Jialong Mai1,2,  Xiaofen Xing1*,  Yawei Li2,  Weidong Chen3 Zhipeng Li1,  Jingyuan Xing1,  Xiangmin Xu1 1South China University of Technology, China 2MiniMax, China  3The Chinese University of Hong Kong, China *Corresponding author 202320111090@mail.scut.edu.cn, xfxing@scut.edu.cn
@@ 1 Introduction
Emotions play a crucial role in human communication. Speech Emotion Recognition (SER) is an important tool for providing user emotional information to intelligent systems (Gross and Muñoz, 1995; Gross et al., 2019). It has wide applications in smart robots, automated call centers, and remote education (Li et al., 2019; Akçay and Oğuz, 2020). Emotion recognition in conversation (ERC) extends SER to enhance emotion recognition capabilities based on conversation history.
@@
Traditional emotion recognition relies on architectures with pretrained representations and well-designed downstream classification networks (Mai et al., ; Chen et al., 2022c). Recently, speech large language model (SLLM), which incorporates speech understanding into language models, has significantly pushed the boundaries of various speech-related tasks. In particular, their potential for SER has garnered increasing attention (Yang et al., 2024a; Bukhari et al., 2024).
@@
Additionally, even training models with longer context windows to alleviate capacity limitations (Ding et al., 2025; Yang et al., 2024a), conventional attention mechanisms (Vaswani et al., 2023) suffer from quadratic computational complexity with respect to input sequence length, rendering them inefficient for long audio inputs.
@@
To address the emotional long dialogue audio processing problem from its source, inspired by Temporary LoRA (Wang et al., 2024), we introduce the Dynamic Parameter Memory (DPM) mechanism. DPM enables models to process unlimited-length audio sequences with linear computational complexity under limited context window conditions. Specifically, we designed an innovative method to "memorize" audio information sentence by sentence into a temporary LoRA module. To support DPM implementation, we first trained an emotion SLLM as the base architecture for DPM inference.
To verify DPM’s effectiveness, we conducted ERC experiments on the IEMOCAP dataset (Busso et al., 2008). Experimental results demonstrate that, within the same emotion-aware SLLM framework, the simple integration of the DPM inference mechanism substantially improves the performance of emotion recognition on long audio sequences, resulting in a 6.56% improvement in weighted accuracy (WA). When processing audio sequences of various lengths, our DPM also brings a 2.23% improvement in WA, demonstrating its robustness across different input durations.
Furthermore, our proposed emotion SLLM itself outperforms traditional classifier approaches, demonstrating the unique advantages of autoregressive structures in audio sequence understanding. In addition, comparative experimental results show that our method has reached state-of-the-art (SOTA) performance. The code is publicly available at our official repository [^1] https://github.com/yongaifadian1/Dynamic-Parameter-Memory_DPM.
@@
  • We designed and trained an effective emotion SLLM that comprehensively outperforms traditional classifier approaches, establishing a solid foundation for application of the DPM mechanism.
  • We propose the Dynamic Parameter Memory (DPM) mechanism, fundamentally solving the difficulty of SLLM processing long emotional speech in conversation, providing a general method with broad application prospects for various audio understanding tasks.
  • We validated DPM’s significant advantages in processing long audio sequences through ERC tasks, proving this method is equally effective across audio of various lengths, with comparative experiments demonstrating its SOTA performance.
@@ 2.1 Downstream Networks for SER
For sentence-level emotion classification tasks, Chen et al. (Chen et al., 2022d) attached trainable downstream networks to frozen wav2vec pretrained models to dynamically capture emotional information in audio; Li et al. (Li et al., 2024) connected improved Transformer downstream networks to frozen HuBERT pretrained models to capture emotional information at different time scales; Mai et al. (Mai et al., ) attached Dropformer networks that dynamically discard noise to frozen WavLM pretrained representations. For ERC, Li et al. (Li et al., 2025) connected improved LSTM networks capable of explicitly capturing conversation history to frozen CLAP models (Wu et al., 2024). Dutta et al. (Dutta and Ganapathy, 2025) attached carefully designed trainable recurrent neural networks to frozen CARE models (Dutta and Ganapathy, 2024).
These approaches greatly leverage and transfer the general knowledge from pretrained representations, making them effective for SER.
@@ 2.3 Enabling LLM to Process Longer Sequences
Murph et al. proposed a method (Murph et al., 2024) that condenses historical information into fixed-length embeddings, which fundamentally alleviates the difficulty of LLMs processing long audio, but potentially suffers from information loss. Chen et al. introduced Temporary LoRA (Wang et al., 2024), which progressively encodes historical text information into parameter space, effectively addressing the challenges of limited-window LLM in understanding long texts. However, this approach lacks validation in other modalities, particularly in the audio domain where high frame rates always present significant challenges.
3 Methodology
The training process of our emotion SLLM and the inference process of DPM are shown in Figure 2. We first train an emotion SLLM as a foundation model with audio understanding and emotion reasoning capabilities for DPM inference. The implementation details are described below.
@@ 3.1 Emotion SLLM
[…]
@@
[…]
@@ 3.1.1 Audio Autoregressive Training
[…]
$\mathcal{L}_{a}$ $=$ (3)
$-\frac{1}{n_{o}}$ $\sum_{k=1}^{n_{o}}\log P\left(x_{T-n_{o}+k}\,\big{|}\,x_{T-n_{o}-n_{p}:T-n_{o}+k-1}\right)$
[…]
@@
[…]
@@ 3.1.2 Emotion Supervision Training
[…]
@@ 3.2 DPM Inference
In the previous section, we described how we trained our emotion SLLM foundation model with audio understanding and emotion classification abilities. In this section, we explain how to implement unlimited-length sequence understanding with a limited LLM context window using DPM, and perform the final emotion inference.
@@
For instance, during pretraining, the emotion SLLM may have developed a relatively neutral understanding of audio. During DPM inference, if the first sentence transcription is "I’ve been waiting for you for an hour," the emotion SLLM might predict that the second sentence would be "I feel tired" based on its general knowledge of audio understanding and emotional patterns. However, if this conversation actually involves an argument, the ground truth for the second sentence might be "I feel angry." This presents a critical challenge: how to "remember" the content of the second sentence, and how to propagate the emotion of anger from this long sample to the inference of subsequent sentences to aid in the final emotion judgment.
@@
$\mathcal{L}_{\text{t}}$ $=$ (9)
$-\frac{1}{m}$ $\sum_{k=1}^{m}\log P\left(x_{T-m+k}\,\big{|}\,x_{T-m-n_{r}}:x_{T-m+k-1}\right)$
@@
[…]
@@
[…]
11:        $\text{loss}\leftarrow\text{loss\_fn}(&\text{pred\_tokens},\\ &\text{sentence.tokens})\end{aligned}$
@@ 4.1 Datasets
The IEMOCAP dataset (Busso et al., 2008) consists of 151 video recordings split into 5 sessions. Keeping in line with previous works (Li et al., 2025; Kieu et al., 2025), we perform a four-class classification task where “angry”, “happy”, “sad”, and “neutral” categories are considered (with excited and happy categories merged). As a result, a total of 5531 utterances spanning four emotion categories are collected. For data partitioning, we adopt the popular “LOSO” (Leave-One-Session-Out) strategy.
@@ 4.2 Experiment Setup
To test the basic capabilities of the CosyVoice2 tokenizer and verify the necessity of the autoregressive structure, we set up comparison experiments with traditional classifier approaches. Specifically, we input the speech tokenizer’s output codes into the emotion SLLM tokenizer, obtain a 4096-dimensional output, and feed it into a downstream network consisting of a trainable Transformer encoder and classifier for classification. The learning rate for the emotion SLLM is 5e-5, and 1e-4 for the classifier. Both are trained for 20 epochs with a batch size of 1. In DPM inference, the temporary LoRA’s learning rate is 5e-5. All experiments are conducted on one NVIDIA A800 GPU.
@@ 4.3.1 Ablation Study
Method* WA (%) UA (%) WF1 (%)
SLLM-DPM* 79.38 79.62 79.34
SLLM* 72.82 73.34 73.58
Classifier* 70.96 70.51 70.64
Method WA (%) UA (%) WF1 (%)
SLLM-DPM 75.38 74.67 74.99
SLLM 73.15 71.98 73.12
Classifier 66.19 63.40 65.22
Table 1: Methods marked with * indicate experiments on complete dialogue samples, others are experiments without length restrictions, covering lengths from single sentences to full dialogues.
Method Year WA (%) UA (%) WF1 (%)
ESA-CRF (Chen et al., 2022a) 2022 73.17 74.47 -
MER-HAN (Zhang et al., 2023) 2023 73.33 74.20 73.66
Mi-CGA (Kieu et al., 2025) 2025 - - 73.77
MER-TL (Padi et al., 2022) 2022 75.76 76.07 -
GatedxLSTM (Li et al., 2025) 2025 76.34 - 75.97
MERITS-L (Dutta and Ganapathy, 2025) 2025 - - 77.95
DPM (ours) 2025 79.38 79.62 79.34
Table 2: Comparison with recent models for ERC.
As shown in Table 1, we present the results of the ablation study, the upper part shows the emotion recognition performance on complete dialogue samples, while the lower part shows the performance on emotion recognition without length constraints, covering lengths from single sentences to full dialogues, aiming to provide a more comprehensive evaluation of DPM’s performance when handling different lengths.
SLLM-DPM represents DPM inference on emotion SLLM, SLLM represents performing a one-time autoregressive inference on emotion SLLM without using DPM, and Classifier represents using a classifier after the speech tokenizer instead of LLM for training and inference. For complete dialogue emotion experiments, we show emotion recognition results only from complete dialogue audio (length ranging from 3-5 minutes). We observe that for understanding emotions in long audio sequences, simply changing the inference method to DPM significantly outperforms direct inference with emotion SLLM (inputting the complete audio sample at once and then inferring the emotion identifier), with a 6.56% improvement in WA. Compared to the classifier baseline, it is 8.42% higher.
This clearly demonstrates the advantage of DPM inference for emotion understanding in long audio sequences. In contrast to processing the entire sequence at once (akin to human "skim-reading"), which can lead to missing critical emotional information, DPM incrementally processes and retains information block by block, resembling a "line-by-line" reading strategy. This enables more precise capture of emotional information in long audio sequences.
Even when extending the duration range of the test audio to cover all length ranges, DPM inference still maintains a significant advantage, with WA exceeding direct emotion SLLM inference by 2.23%, demonstrating the general applicability of DPM inference.
@@
When using emotion SLLM for direct inference without DPM, our method outperformed traditional classifier approaches in both settings, indicating the unique advantage of the autoregressive structure over classifiers in understanding emotions in audio sequences.
It’s worth noting that since DPM inference computation is proportional to the number of sentences, DPM inference achieves linear computational complexity. When audio sequence length increases, inference computational complexity grows linearly.

### v2

@@ Abstract
Recent research has focused on applying speech large language model (SLLM) to improve speech emotion recognition (SER). However, the inherently high frame rate in speech modality severely limits the signal processing and understanding capabilities of SLLM. For example, a SLLM with a 4K context window can only process 80 seconds of audio at 50Hz feature sampling rate before reaching its capacity limit. Input token compression methods used in SLLM overlook the continuity and inertia of emotions across multiple conversation turns. This paper proposes a Dynamic Parameter Memory (DPM) mechanism with contextual semantics and sentence-level emotion encoding, enabling processing of unlimited-length audio with limited context windows in SLLM. Specifically, DPM progressively encodes sentence-level information and emotions into a temporary LoRA module during inference to effectively ”memorize” the contextual information. We trained an emotion SLLM as a backbone and incorporated our DPM into inference for emotion recognition in conversation (ERC). Experimental results on the IEMOCAP and MELD datasets show that DPM significantly improves the emotion recognition capabilities of SLLM when processing long audio sequences, achieving state-of-the-art performance.
@@ 1 Introduction
Emotions play a crucial role in human communication. Speech Emotion Recognition (SER) is an important tool for providing user emotional information to intelligent systems (Gross & Muñoz, 1995; Gross et al., 2019). It has wide applications in smart robots, automated call centers, and remote education (Li et al., 2019; Akçay & Oğuz, 2020). Emotion recognition in conversation (ERC) extends SER to enhance emotion recognition capabilities based on conversation history.
@@
Traditional emotion recognition relies on architectures with pretrained representations and well-designed downstream classification networks (Mai et al., 2024; Chen et al., 2022c). Recently, speech large language model (SLLM), which incorporates speech understanding into language models, has significantly pushed the boundaries of various speech-related tasks. In particular, their potential for SER has garnered increasing attention (Yang et al., 2024a; Bukhari et al., 2024).
@@
Additionally, even training models with longer context windows to alleviate capacity limitations (Ding et al., 2025; Yang et al., 2024a), conventional attention mechanisms (Vaswani et al., 2017) suffer from quadratic computational complexity with respect to input sequence length, rendering them inefficient for long audio inputs.
@@
We introduce the Dynamic Parameter Memory (DPM), a novel inference mechanism designed to process unlimited-length emotional audio dialogues within a limited context window. While the concept of updating parameters at inference time, as explored by Temporary LoRA (Wang et al., 2024) for text, provides a promising direction, its direct application to speech is non-trivial due to the high temporal density and unique structure of emotional expression in audio. To bridge this gap, our core contribution is a two-fold approach: first, we develop an emotion-aware SLLM specifically pre-trained to comprehend emotional nuances in audio sequences. Second, we build the DPM mechanism upon this foundation, which iteratively distills sentence-level audio information and emotional context into a temporary LoRA module. This allows our method to maintain a dynamically updated ”memory” of the conversation, achieving a deep, context-aware understanding of long dialogues.
To verify DPM’s effectiveness, we conduct extensive ERC experiments on two standard benchmarks: IEMOCAP (Busso et al., 2008) and MELD (Poria et al., 2018). Experimental results consistently demonstrate that, within the same emotion-aware SLLM framework, the simple integration of the DPM inference mechanism substantially improves emotion recognition performance, particularly on long audio sequences. Furthermore, our method shows robust gains across samples of various lengths, confirming its general applicability.
Furthermore, our proposed emotion SLLM itself outperforms traditional classifier approaches, demonstrating the unique advantages of autoregressive structures in audio sequence understanding. Comparative experiments show that our DPM-enhanced method achieves state-of-the-art (SOTA) performance on both datasets.
@@
  • We design and train an effective emotion SLLM that comprehensively outperforms traditional classifier approaches, establishing a solid foundation for the application of the DPM mechanism.
  • We propose the Dynamic Parameter Memory (DPM) mechanism, a novel inference strategy that fundamentally addresses the challenge of processing long emotional speech in SLLM, enabling the handling of unlimited-length audio with linear computational complexity under a limited context window.
  • We conduct extensive experiments on the IEMOCAP and MELD datasets, demonstrating that our DPM-enhanced SLLM achieves new state-of-the-art results on both benchmarks.
  • We provide an in-depth analysis of DPM’s stepping strategies, revealing a critical trade-off between performance and efficiency and reinforcing the value of using semantic units for memory updates.
@@ 2.3 Enabling LLM to Process Longer Sequences
Murph et al. proposed a method (Murph et al., 2024) that condenses historical information into fixed-length embeddings, which fundamentally alleviates the difficulty of LLMs processing long audio, but potentially suffers from information loss. Wang et al. introduced Temporary LoRA (Wang et al., 2024), which progressively encodes historical text information into parameter space, effectively addressing the challenges of limited-window LLM in understanding long texts. However, this approach lacks validation in other modalities, particularly in the audio domain where high frame rates always present significant challenges.
3 Methodology
The training process of our emotion SLLM and the inference process of DPM are shown in Figure 2. We first train an emotion SLLM as a foundation model with audio understanding and emotion reasoning capabilities for DPM inference. The implementation details are described below, using our setup for the IEMOCAP dataset (four-class classification) as the primary example.
@@ 3.1 Emotion SLLM
To enable the LLM to process these discrete audio codes, we first expand the vocabulary size of the LLM tokenizer to accommodate the speech tokenizer’s codebook size, represented as $\langle\text{audio}\_{i}\rangle$. We also add a token representing the end of the current sentence, denoted as $\langle\text{audio\_end}\rangle$, and four emotion identifiers represented as $\langle\text{emo\_hap}\rangle$, $\langle\text{emo\_sad}\rangle$, $\langle\text{emo\_ang}\rangle$, and $\langle\text{emo\_neu}\rangle$.
@@
We use LoRA (Hu et al., 2021) to train the SLLM. Before training, we preprocess the training data. For a long audio sample, we first convert the audio into discrete codes using a speech tokenizer. Then, we insert $\langle\text{audio\_end}\rangle$ at the end of each sentence, followed by the corresponding emotion identifier. For example, for a 5-minute audio consisting of 50 sentences, we would insert 50 $\langle\text{audio\_end}\rangle$ tokens and 50 corresponding emotion identifiers at the end of each sentence.
@@ 3.1.1 Audio Autoregressive Training
We train the model’s next token prediction to ensure its understanding of audio. For each sentence, we take $n_{o}$ tokens before $\langle\text{audio\_end}\rangle$ (including $\langle\text{audio\_end}\rangle$) as the training target for autoregressive prediction. We take $n_{p}$ tokens before these as the prefix for predicting the next $n_{o}$ tokens, using teacher forcing for autoregressive training. The loss for these $n_{o}$ tokens is denoted as $\mathcal{L}_{a}$.
$\mathcal{L}_{a}=-\frac{1}{n_{o}}$ $\sum_{k=1}^{n_{o}}\log P\left(x_{T-n_{o}+k}\,\big|\,x_{T-n_{o}-n_{p}:T-n_{o}+k-1}\right)$ (3)
where $T$ represents the position of $\langle\text{audio\_end}\rangle$, $x$ is a sequence of audio token after preprocessing.
@@
Through audio autoregressive training, the emotion SLLM learns to autoregressively predict the next audio token when receiving an audio token and predict $\langle\text{audio\_end}\rangle$ at appropriate times, gaining audio understanding ability.
@@ 3.1.2 Emotion Supervision Training
We also train the model’s emotion classification ability. For each sentence ending, we take the preceding $n_{q}$ tokens (including $\langle\text{audio\_end}\rangle$) as the prefix for predicting the emotion identifier. Notably, for the emotion SLLM’s output, we constrain the logits from the original size of $|V{\text{text}}|+|\langle\text{audio}\_{i}\rangle|+1+4$ to just 4 (position of emotion identifiers) to increase the stability of emotion prediction, constrained logits denoted as $\hat{p}$. We then calculate the cross-entropy loss $\mathcal{L}_{e}$ using the constrained logits and the true emotion identifier.
@@ 3.2 DPM Inference
For instance, during pretraining, the emotion SLLM may have developed a relatively neutral understanding of audio. During DPM inference, if the first sentence transcription is ”I’ve been waiting for you for an hour,” the emotion SLLM might predict that the second sentence would be ”I feel tired” based on its general knowledge of audio understanding and emotional patterns. However, if this conversation actually involves an argument, the ground truth for the second sentence might be ”I feel angry.” This presents a critical challenge: how to ”remember” the content of the second sentence, and how to propagate the emotion of anger from this long sample to the inference of subsequent sentences to aid in the final emotion judgment.
@@
$\mathcal{L}_{\text{t}}=-\frac{1}{m}$ $\sum_{k=1}^{m}\log P\left(x_{T-m+k}\,\big|\,x_{T-m-n_{r}}:x_{T-m+k-1}\right)$ (9)
@@
3:  $\text{temp\_lora}\leftarrow\text{SLLM.create\_templora}()$
@@
9:   $\text{prefix}\leftarrow\text{sample.tokens}[T-n_{r}:T]$
10:   $\text{pred\_tokens}\leftarrow\text{SLLM.infer}(\text{prefix})$
11:   $\begin{aligned} \text{loss}\leftarrow\text{loss\_fn}(&\text{pred\_tokens},\text{sentence.tokens})\end{aligned}$
@@ 4.1 Datasets
IEMOCAP
The IEMOCAP dataset (Busso et al., 2008) consists of 151 video recordings split into 5 sessions. Keeping in line with previous works (Li et al., 2025; Kieu et al., 2025), we perform a four-class classification task where “angry”, “happy”, “sad”, and “neutral” categories are considered (with the excited category merged into happy). As a result, a total of 5531 utterances spanning four emotion categories are collected. For data partitioning, we adopt the popular “LOSO” (Leave-One-Session-Out) strategy.
MELD
The MELD (Poria et al., 2018) is a multi-speaker conversational dataset derived from the TV series Friends. It comprises 1,433 dialogues from 407 different speakers. The dataset is officially partitioned into training (1,039 dialogues), validation (114 dialogues), and test (280 dialogues) sets. Each utterance is annotated with one of seven emotion categories: “angry”, “disgust”, “sad”, “joy”, “neutral”, “surprise”, and “fear”. We adhere to the official data splits for our experiments.
@@ 4.2 Experiment Setup
To establish a strong baseline and isolate the gains from our SLLM’s autoregressive structure, we compared our model against a traditional classifier-based approach. This baseline model processes the speech tokenizer’s output codes through a trainable Transformer encoder followed by a classification head. For training, the learning rate for the emotion SLLM was set to 5e-5, while the classifier baseline used 1e-4. For our DPM inference, the temporary LoRA module was updated with a learning rate of 5e-5. All models were trained for 20 epochs, and all experiments were conducted on a single NVIDIA A800 GPU.
The prefix length for DPM inference (the number of tokens preceding the final emotion token used for prediction) was tuned as a hyperparameter for each dataset. We set it to 1024 for IEMOCAP and 256 for MELD. This choice aligns with the datasets’ characteristics: IEMOCAP features significantly longer dialogues (avg. 64.96 utterances) compared to MELD (avg. 9.80 utterances). For the lengthy and complex conversations in IEMOCAP, a larger prefix provides a more robust ”recent context” to anchor the extensive parametric memory accumulated by the temporary LoRA. Conversely, the more concise dialogues in MELD require a shorter prefix for an effective and efficient final prediction.
@@ 4.3.1 Ablation Study
Table 1: Ablation study on IEMOCAP and MELD. We compare our full model (SLLM-DPM) against its variants under two settings: (1) using only complete dialogue samples and (2) using samples of all lengths.
Method IEMOCAP MELD
WA (%) UA (%) WF1 (%) WF1 (%)
Setting 1: On complete dialogue samples
SLLM-DPM 79.38 79.62 79.34 51.22
SLLM 72.82 73.34 73.58 47.90
Classifier 70.96 70.51 70.64 44.78
Setting 2: On samples of all lengths
SLLM-DPM 75.38 74.67 74.99 52.82
SLLM 73.15 71.98 73.12 52.53
Classifier 66.19 63.40 65.22 46.84
As shown in Table 1, we present the results of the ablation study. The upper part shows the emotion recognition performance on complete dialogue samples, while the lower part shows the performance on emotion recognition without length constraints, covering lengths from single sentences to full dialogues, aiming to provide a more comprehensive evaluation of DPM’s performance when handling different lengths.
SLLM-DPM represents DPM inference on emotion SLLM, SLLM represents performing a one-time autoregressive inference on emotion SLLM without using DPM, and Classifier represents using a classifier after the speech tokenizer instead of LLM for training and inference. For complete dialogue emotion experiments, we show emotion recognition results only from complete dialogue audio. We observe that for understanding emotions in long audio sequences, simply changing the inference method to DPM significantly outperforms direct inference with emotion SLLM (inputting the complete audio sample at once and then inferring the emotion identifier), with a 6.56% improvement in WA on IEMOCAP. A similar trend is observed on the MELD dataset, where DPM achieves a 3.32% gain in WF1. Compared to the classifier baseline, our full model shows an 8.42% improvement in WA on IEMOCAP and a 6.44% improvement in WF1 on MELD.
This clearly demonstrates the advantage of DPM inference for emotion understanding in long audio sequences. In contrast to processing the entire sequence at once (similar to human ”skim-reading”), which can lead to missing critical emotional information, DPM incrementally processes and retains information block by block, resembling a ”line-by-line” reading strategy. This enables more precise capture of emotional information in long audio sequences.
Even when extending the duration range of the test audio to cover all length ranges, DPM inference still maintains a significant advantage, with WA exceeding direct emotion SLLM inference by 2.23% on IEMOCAP, and showing a gain on MELD as well. This demonstrates the general applicability of DPM inference.
@@
When using emotion SLLM for direct inference without DPM, our method outperformed traditional classifier approaches in both settings and across both datasets, indicating the unique advantage of the autoregressive structure over classifiers in understanding emotions in audio sequences.
It’s worth noting that since DPM inference computation is proportional to the number of sentences, DPM inference achieves linear computational complexity. When audio sequence length increases, inference computational complexity grows linearly. In addition to these main results, we conducted further experiments to analyze the DPM stepping strategies and the impact of context window size on performance. These detailed analyses are presented in the Appendix.
Table 2: Comparison with recent models for ERC on IEMOCAP and MELD datasets. The best results are in bold.
Method IEMOCAP MELD
WA (%) UA (%) WF1 (%) WF1 (%)
ESA-CRF (Chen et al., 2022a) 73.17 74.47 - -
MER-HAN (Zhang et al., 2023) 73.33 74.20 73.66 40.50
Mi-CGA (Kieu et al., 2025) - - 73.77 -
MER-TL (Padi et al., 2022) 75.76 76.07 - -
GatedxLSTM (Li et al., 2025) 76.34 - 75.97 -
MERITS-L (Dutta & Ganapathy, 2025) - - 77.95 49.32
DPM (ours) 79.38 79.62 79.34 51.22